# Final-Model-Only Data Attribution with a Unifying View of Gradient-Based Methods

### v1

@@ Abstract
Training data attribution (TDA) is the task of attributing model behavior to elements in the training data. This paper draws attention to the common setting where one has access only to the final trained model, and not the training algorithm or intermediate information from training. To serve as a gold standard for TDA in this “final-model-only” setting, we propose further training, with appropriate adjustment and averaging, to measure the sensitivity of the given model to training instances. We then unify existing gradient-based methods for TDA by showing that they all approximate the further training gold standard in different ways. We investigate empirically the quality of these gradient-based approximations to further training, for tabular, image, and text datasets and models. We find that the approximation quality of first-order methods is sometimes high but decays with the amount of further training. In contrast, the approximations given by influence function methods are more stable but surprisingly lower in quality.
@@ 1 Introduction
Training data attribution (TDA, or sometimes simply “data attribution”) refers to the attribution or explanation of ML model behavior in terms of its training data. Existing methods for TDA can be broadly categorized into re-training-based approaches and gradient-based approaches. The former perform attribution by re-training on different subsets of training data and observing the corresponding effects (Ghorbani and Zou, 2019; Ghorbani et al., 2020; Feldman and Zhang, 2020; Kwon and Zou, 2022; Ilyas et al., 2022; Lin et al., 2022; Wang and Jia, 2023). The latter are in turn divided into those that track the effects of training instances throughout gradient-based training (Hara et al., 2019; Pruthi et al., 2020; Chen et al., 2021; Bae et al., 2024) versus those that are applied only at the end (Koh and Liang, 2017; Yeh et al., 2018; Koh et al., 2019; Barshan et al., 2020; Basu et al., 2020; Guo et al., 2021; Schioppa et al., 2022; Park et al., 2023; Kwon et al., 2024). We refer the reader to the survey by Hammoudeh and Lowd (2024) for a deeper look at these categories.
@@
In this paper, rather than proposing another TDA method, we take a reflective approach to the literature. First, we draw attention to the fact that there exist multiple problem settings for TDA, alongside the multiple categories of methods. These problem settings differ in the level of access assumed. In particular, we focus on what we call the “final-model-only” (FiMO) setting in which we have access only to the final trained model, and not the algorithm used to train the model or intermediate information from training (for example checkpoints). The FiMO setting is motivated by the common scenario in which TDA is performed by a different party than the one who developed the model. This is the case for example for models downloaded from platforms such as HuggingFace.
@@
We find that since the TDA literature has not clearly differentiated these problem settings, it is also not clear on what should be the ideal “gold standard” for TDA in the FiMO setting. Having a gold standard (as opposed to proxy tasks such as mislabelled example detection) facilitates the development and evaluation of more practical methods. We thus propose further training, starting from the given final model, as a gold standard measure of the sensitivity of the model to training instances. Our proposal adjusts for the effect of further training alone (as distinct from the effects of particular training instances) and accounts for the randomness of typical neural network training algorithms.
@@
Like re-training [^1] Throughout the paper, we reserve the term “re-training” to mean re-training from scratch., further training a model multiple times can be computationally prohibitive. We thus consider approximations based on first- and second-order Taylor expansions of the further training objective. In doing so, we unify several existing gradient-based methods for TDA. These methods include gradient similarity (Charpiat et al., 2019) (a final-checkpoint-only case of TracIn (Pruthi et al., 2020)) and methods based on influence functions (Koh and Liang, 2017; Schioppa et al., 2022; Park et al., 2023; Grosse et al., 2023; Kwon et al., 2024). The direct connection that these gradient-based methods have to our further training objective suggests that they are more suited to the FiMO setting, rather than as approximations to re-training where their effectiveness has been questioned (Basu et al., 2021; Nguyen et al., 2023).
@@
We investigate empirically the quality of the approximations to further training provided by different gradient-based methods. Our experiments span the modalities of tabular, image, and text data. Overall, we find that first-order gradient-based methods can give good initial approximations to further training, but the quality of approximation decays with the amount of further training. In contrast, the approximation quality of influence function methods is more persistent, but somewhat surprisingly, never as high as first-order methods at their peak.
@@
Our contributions can be summarized as follows. 1) We highlight the FiMO setting for TDA (Section 2). 2) We articulate a further training gold standard for TDA in the FiMO setting (Section 3). 3) We show how several gradient-based TDA methods approximate further training, theoretically (Section 4) and numerically (Section 6). We discuss additional implications of our work in Section 7.
@@ 2 Problem Settings for Training Data Attribution
In all problem settings that we consider for training data attribution (TDA), we are given access to the training dataset $\mathcal{D}=\{{\bm{z}}_{i}\}_{i=1}^{n}$ for the model, consisting of $n$ pairs ${\bm{z}}_{i}=({\bm{x}}_{i},y_{i})$ of inputs ${\bm{x}}_{i}\in{\mathcal{X}}$ and targets $y_{i}\in{\mathcal{Y}}$. A model $f({\bm{x}};{\bm{\theta}})$ is a function $f:{\mathcal{X}}\to{\mathcal{F}}$, parameterized by ${\bm{\theta}}\in\mathbb{R}^{p}$, that maps to an output space ${\mathcal{F}}$ (e.g., predicted logits for a classification model). The quality of each model output $f({\bm{x}}_{i};{\bm{\theta}})$ with respect to target $y_{i}$ is measured by a loss function $\ell(f({\bm{x}}_{i};{\bm{\theta}}),y_{i})$; we adopt the more compact notation $L({\bm{z}}_{i};{\bm{\theta}})\triangleq\ell(f({\bm{x}}_{i};{\bm{\theta}}),y_{i})$. We use $A$ to denote the training algorithm or learning algorithm used to train $f({\bm{x}};{\bm{\theta}})$ on dataset $\mathcal{D}$. We view $A$ as a function that takes $\mathcal{D}$ and initial model parameters ${\bm{\theta}}^{(0)}$ (which may come from a pre-trained model) as input and outputs final model parameters ${\bm{\theta}}^{f}=A(\mathcal{D},{\bm{\theta}}^{(0)})$. The algorithm $A$ may be stochastic, for example due to the random ordering of instances in each training epoch.
Given a test instance ${\bm{z}}=({\bm{x}},y)$ and an evaluation function $g({\bm{z}},{\bm{\theta}})$ that depends in general on both $f({\bm{x}};{\bm{\theta}})$ and $y$, the task of TDA is to assign scores $a_{i}$ quantifying the importance of each training instance ${\bm{z}}_{i}$ to the output $g({\bm{z}},{\bm{\theta}})$. We refer to $a_{i}$ interchangeably as an attribution or influence score. Commonly, the evaluation function is the loss on ${\bm{z}}$, $g({\bm{z}},{\bm{\theta}})=L({\bm{z}};{\bm{\theta}})$, or the output of the model alone, $g({\bm{z}},{\bm{\theta}})=f({\bm{x}};{\bm{\theta}})$. One may also consider evaluation functions that sum over test instances, i.e., $\sum_{{\bm{z}}\in\mathcal{D_{\mathrm{test}}}}g({\bm{z}},{\bm{\theta}})$.
@@
    Final Model Only (FiMO): We have access only to the final model $f({\bm{x}};{\bm{\theta}}^{f})$ and not the training algorithm $A$ or intermediate information. The algorithm $A$ may be actually unavailable, or it may be available but too computationally expensive or otherwise undesirable to re-run. This scenario typically occurs when model training and data attribution are performed by different parties.
@@
The three settings also differ in the object whose behavior is being attributed. In the TAA setting, the object is the algorithm $A$ and its sensitivity to different training instances ${\bm{z}}_{i}$. In the CPA setting, it is the training trajectory. The FiMO setting is the most focused on the final model $f({\bm{x}};{\bm{\theta}}^{f})$ that will actually be used.
@@
In the remainder of the paper, we focus on the FiMO setting as it requires the least amount of access and is thus the most widely applicable. It applies for example to open-weights models that are freely downloadable from platforms such as HuggingFace. By extension, the FiMO setting applies in cases where methods for the TAA and CPA settings (i.e., re-training and tracing methods) cannot be used.
Figure 1: Given only a final model with parameters ${\bm{\theta}}^{f}$, the proposed further training gold standard measures the model’s sensitivity to training sample $i$ by conducting a limited amount of further training on the full training set $\mathcal{D}$ as well as leaving out sample $i$ ($\mathcal{D}_{-i}$), resulting in changed parameters ${\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D})$ and ${\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}_{-i})$. The difference in outputs $g({\bm{z}},{\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}_{-i}))-g({\bm{z}},{\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}))$ between the two further trained models indicates the sensitivity to $i$. If training is stochastic, then this process should be repeated to obtain the expected sensitivity.
3 A Further Training Gold Standard for the FiMO Setting
@@
We first consider the question of how TDA should ideally be done in the FiMO setting, i.e., what could serve as a “gold standard” method. While a gold standard method may be very expensive to carry out in many cases, it is nevertheless useful in evaluating and developing approximate TDA methods that are more practical. However, since the TDA literature has not explicitly delineated the FiMO setting to our knowledge, it is also not clear what the gold standard should be.
@@
Returning for a moment to the TAA setting where the training algorithm is available, the natural question to ask there is one of contribution: how much does a training instance ${\bm{z}}_{i}$ contribute to the model through the training process? Accordingly, variants of re-training, designed to estimate these contributions, are accepted as gold standards in the TAA setting. The simplest of these is leave-one-out (LOO) re-training where each instance ${\bm{z}}_{i}$ is left out of the training set in turn to assess its contribution. In the FiMO setting however, it is not clear how one can “go back in time” to determine contributions to the final model. Instead, we change the question to one of sensitivity: how sensitive is the given model to a training instance ${\bm{z}}_{i}$? To measure these sensitivities, we propose variants of further training as a gold standard, in a manner analogous to how re-training is used in the TAA setting.
@@
Further training can be described generically as follows. We start from the given parameters ${\bm{\theta}}^{f}$ and train on a dataset $\mathcal{D}^{\prime}$ using an algorithm $A^{\prime}$, i.e., ${\bm{\theta}}^{f}+\Delta{\bm{\theta}}=A^{\prime}(\mathcal{D}^{\prime},{\bm{\theta}}^{f})$, where $\mathcal{D}^{\prime}$ and $A^{\prime}$ are generally different from $\mathcal{D}$ and $A$. In this work, we restrict attention to LOO further training, corresponding to LOO datasets $\mathcal{D}^{\prime}=\mathcal{D}_{-i}\triangleq\mathcal{D}\setminus\{{\bm{z}}_{i}\}$ as well as $\mathcal{D}^{\prime}=\mathcal{D}$, but more general perturbed datasets $\mathcal{D}^{\prime}$ are possible. As for $A^{\prime}$, our desire is for it to be representative of typical neural network (NN) training. We thus assume that $A^{\prime}$ seeks to minimize the empirical risk $R(\mathcal{D}^{\prime};{\bm{\theta}})\triangleq\sum_{{\bm{z}}_{i}\in\mathcal{D}^{\prime}}L({\bm{z}}_{i};{\bm{\theta}})$:
$\Delta{\bm{\theta}}(\mathcal{D}^{\prime})\approx\operatorname*{arg\,min}_{\Delta\theta}\>R(\mathcal{D}^{\prime};{\bm{\theta}}^{f}+\Delta{\bm{\theta}}),$ (1)
@@
where the notation ${\bm{\theta}}^{f}+\Delta{\bm{\theta}}$ is meant to indicate that the optimization is initialized at ${\bm{\theta}}^{f}$ ($\Delta{\bm{\theta}}={\bm{0}}$), and $\Delta{\bm{\theta}}(\mathcal{D}^{\prime})$ is the change in parameters resulting from applying $A^{\prime}$ to $(\mathcal{D}^{\prime},{\bm{\theta}}^{f})$.
@@
With further training as a building block, we now address two issues inherent to typical NN training. [^2] For convex training objectives, these issues are not much of a concern. Moreover, further training and re-training are essentially equivalent (please see Appendix A).
Non-convergence:
@@
The “final” parameters ${\bm{\theta}}^{f}$ are often not a stationary point of the empirical risk. Hence, as shown in Figure 1, further training on the original training set $\mathcal{D}^{\prime}=\mathcal{D}$ generally yields a non-zero change $\Delta{\bm{\theta}}(\mathcal{D})$. This parameter shift is reflected in the evaluation function as $g({\bm{z}},{\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}))$ and can be interpreted as the effect due to further training alone. Similarly, $\Delta{\bm{\theta}}(\mathcal{D}_{-i})$ and $g({\bm{z}},{\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}_{-i}))$ in Figure 1 are the effects of further training without instance ${\bm{z}}_{i}$. The difference in outputs,
@@
Stochasticity of training:
The training algorithm $A^{\prime}$ also depends on random elements, notably the order of instances in each training epoch. Denoting these random elements as $\xi$, the change in parameters becomes a function of $\xi$, ${\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}^{\prime},\xi)=A^{\prime}(\mathcal{D}^{\prime},{\bm{\theta}}^{f},\xi)$, as does the evaluation function $g({\bm{z}},{\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}^{\prime},\xi))$. We view $\xi$ as a nuisance parameter, an artifact of the stochastic nature of modern training algorithms. Thus, we would ideally like to take an expectation over $\xi$. Applying this to the difference in (2) yields

### v2

@@ Abstract
Training data attribution (TDA) is concerned with understanding model behavior in terms of the training data. This paper draws attention to the common setting where one has access only to the final trained model, and not the training algorithm or intermediate information from training. We reframe the problem in this “final-model-only” setting as one of measuring sensitivity of the model to training instances. To operationalize this reframing, we propose further training, with appropriate adjustment and averaging, as a gold standard method to measure sensitivity. We then unify existing gradient-based methods for TDA by showing that they all approximate the further training gold standard in different ways. We investigate empirically the quality of these gradient-based approximations to further training, for tabular, image, and text datasets and models. We find that the approximation quality of first-order methods is sometimes high but decays with the amount of further training. In contrast, the approximations given by influence function methods are more stable but surprisingly lower in quality.
@@ 1 Introduction
Training data attribution (TDA, or sometimes simply “data attribution”) refers to the attribution or explanation of ML model behavior in terms of its training data. Existing methods for TDA fall into several categories: re-training-based approaches (ghorbani2019datashapley; ghorbani202020distributional; feldman2020memorize; kwon2022beta; ilyas2022datamodels; lin2022measuring; wang2023banzhaf), gradient-based approaches applied throughout training (hara2019cleansing; pruthi2020tracin; chen2021hydra; bae2024unrolled), and gradient approaches applied only at the end (koh2017inffunc; yeh2018representer; koh2019group; barshan2020relatif; basu2020second; guo2021fastif; schioppa2022scalingarnoldi; park2023trak; kwon2024datainf). We refer to the survey by hammoudeh2024tdasurvey for a deeper look at these categories.
@@
In this paper, rather than proposing another TDA method, we take a more reflective approach. First, we draw attention to the fact that there exist multiple problem settings for TDA, alongside the multiple categories of methods. These problem settings differ in the level of access assumed. In particular, we focus on what we call the “final-model-only” (FiMO) setting in which we have access only to the final trained model, and not the algorithm used to train the model or intermediate information from training (for example checkpoints). The FiMO setting is motivated by the common scenario in which TDA is performed by a different party than the one who developed the model. Models published on platforms such as HuggingFace have now made this scenario ubiquitous.
@@
We find that since the TDA literature has not clearly differentiated these problem settings, it is also not clear on what should be the goal for TDA in the FiMO setting, and accordingly, what could be an ideal “gold standard” method. Having such a goal and gold standard (as opposed to proxy tasks such as mislabelled example detection) facilitates the development and evaluation of more practical methods. We thus reframe the problem in the FiMO setting as one of quantifying sensitivity of the model to training instances. We call the reframed problem “FiMODA” (DA for data attribution). We then propose further training, starting from the given final model, as a gold standard measure of sensitivity. Notably, our proposal adjusts for the effect of further training non-converged models and accounts for the randomness of neural network training algorithms. These refinements have consequences as we show in our experiments.
@@
Like re-training [^1] Throughout the paper, we reserve the term “re-training” to mean re-training from scratch., further training a model multiple times can be computationally prohibitive. We thus consider approximations based on first- and second-order Taylor expansions of the further training objective. In doing so, we unify several existing gradient-based methods for TDA. These methods include gradient similarity (charpiat2019similarity) (a final-checkpoint-only case of TracIn (pruthi2020tracin)) and methods based on influence functions (koh2017inffunc; schioppa2022scalingarnoldi; park2023trak; grosse2023studyingllmgen; kwon2024datainf). Given that these gradient-based methods approximate further training as we show, we submit that they are more suited to the FiMO setting, rather than as approximations to re-training where their effectiveness has been questioned (basu2021inffragile; nguyen2023bayesian). Our derivation of gradient-based methods also does not make assumptions of convexity or stationarity that are common in the TDA literature, and we obtain generalized influence function expressions as a result.
@@
We investigate empirically the quality of the approximations to further training provided by different gradient-based methods. Our experiments span the modalities of tabular, image, and text data. Overall, we find that first-order gradient-based methods can give good initial approximations to further training, but the quality of approximation decays with the amount of further training. In contrast, the approximation quality of influence function methods is more persistent, but somewhat surprisingly, never as high as first-order methods at their peak. We provide code to help reproduce our experiments at https://github.com/IBM/fimoda.
Table 1: Comparison with other retrospective works on TDA plus koh2017inffunc (FT = Further training)
Category Attribute Ours koh2017inffunc basu2021inffragile nguyen2023bayesian zhang2022rethinking bae2022infanswer schioppa2023theoretical
Setting FiMO (explicit)
FT refinements Training randomness
Non-convergence
Derivations Non-convexity/stationarity
Generalized influence functions
Scope # TDA methods evaluated 8 2 2 3 2 2 2
@@
Our contributions are summarized as follows and in Table 1: 1) We highlight the FiMO setting for TDA (Section 2). 2) We reframe the TDA problem in the FiMO setting as one of quantifying sensitivity. We articulate and refine a further training gold standard for this reframed problem (called FiMODA, Section 3). 3) We show how several gradient-based TDA methods approximate further training, theoretically (Section 4) and numerically (Section 6).
@@ 2 Problem Settings for Training Data Attribution
In all problem settings that we consider for TDA, we are given access to the training dataset $\mathcal{D}=\{{\bm{z}}_{i}\}_{i=1}^{n}$ for the model, consisting of $n$ pairs ${\bm{z}}_{i}=({\bm{x}}_{i},y_{i})$ of inputs ${\bm{x}}_{i}\in{\mathcal{X}}$ and targets $y_{i}\in{\mathcal{Y}}$. A model $f({\bm{x}};{\bm{\theta}})$ is a function $f:{\mathcal{X}}\to{\mathcal{F}}$, parameterized by ${\bm{\theta}}\in\mathbb{R}^{p}$, that maps to an output space $\mathcal{F}$} (e.g., predicted logits for a classification model). The quality of each model output $f({\bm{x}}_{i};{\bm{\theta}})$ with respect to target $y_{i}$ is measured by a loss function $\ell(f({\bm{x}}_{i};{\bm{\theta}}),y_{i})$; we adopt the more compact notation $L({\bm{z}}_{i};{\bm{\theta}})\triangleq\ell(f({\bm{x}}_{i};{\bm{\theta}}),y_{i})$. We use $A$ to denote the training algorithm used to train $f({\bm{x}};{\bm{\theta}})$ on dataset $\mathcal{D}$. We view $A$ as a function that takes $\mathcal{D}$ and initial model parameters ${\bm{\theta}}^{(0)}$ (which may come from a pre-trained model) as input and outputs final model parameters ${\bm{\theta}}^{f}=A(\mathcal{D},{\bm{\theta}}^{(0)})$.
Given a test instance ${\bm{z}}=({\bm{x}},y)$ and an evaluation function $g({\bm{z}},{\bm{\theta}})$ that depends in general on both $f({\bm{x}};{\bm{\theta}})$ and $y$, the task of TDA is to assign scores $a_{i}$ quantifying the importance of each training instance ${\bm{z}}_{i}$ to the output $g({\bm{z}},{\bm{\theta}})$. We refer to $a_{i}$ interchangeably as an attribution or influence score. Commonly, the evaluation function is the loss on $\bm{z}$}, $g({\bm{z}},{\bm{\theta}})=L({\bm{z}};{\bm{\theta}})$, or the output of the model alone, $g({\bm{z}},{\bm{\theta}})=f({\bm{x}};{\bm{\theta}})$. One may also consider evaluation functions that sum over test instances $\bm{z}$}.
@@
    Final Model Only (FiMO): We have access only to the final model $f({\bm{x}};{\bm{\theta}}^{f})$ and not the training algorithm $A$ or intermediate information. The algorithm $A$ may actually be unavailable, or it may be available but too computationally expensive or otherwise undesirable to re-run. This scenario typically occurs when model training and data attribution are performed by different parties.
@@
The three settings also differ in the object being attributed. In the TAA setting, it is the algorithm $A$ and its sensitivity to different training instances ${\bm{z}}_{i}$. In the CPA setting, it is the training trajectory. The FiMO setting is the most focused on the final model $f({\bm{x}};{\bm{\theta}}^{f})$ that will actually be used.
@@
In the remainder of the paper, we focus on the FiMO setting as it requires the least amount of access and is thus the most widely applicable. It applies for example to open-weights models that are freely downloadable from platforms such as HuggingFace. By extension, the FiMO setting applies in cases where methods for the TAA and CPA settings (i.e., re-training and tracing methods) cannot be used. [^2] Naturally, if the increased access of the TAA and CPA settings is available, then methods that take advantage of this can do more than is possible in the FiMO setting. We will use the term “FiMODA” to refer to TDA in the FiMO setting and to distinguish it from the standard TDA problem.
3 A Further Training Gold Standard for Final-Model-Only Data Attribution
@@
We first consider the question of how FiMODA should ideally be done, i.e., what should be the goal and what could serve as a “gold standard” method, where the gold standard also respects the FiMO constraint. While a gold standard method may be very expensive to carry out in many cases, it is nevertheless useful in evaluating and developing approximate FiMODA methods that are more practical. However, since the TDA literature has not explicitly delineated the FiMO setting to our knowledge, it is also not clear what the goal and gold standard should be.
Figure 1: Given only a final model with parameters ${\bm{\theta}}^{f}$, the proposed further training gold standard measures the model’s sensitivity to training sample $i$ by further training on the full training set $\mathcal{D}$ as well as leaving out sample $i$ ($\mathcal{D}_{-i}$), resulting in changed parameters ${\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D})$ and ${\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}_{-i})$. The difference in outputs $g({\bm{z}},{\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}_{-i}))-g({\bm{z}},{\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}))$ indicates the sensitivity to $i$. For stochastic training, this process is repeated to obtain the expected sensitivity.
@@
Returning for a moment to the TAA setting, the natural question to ask there is one of contribution: how much does a training instance ${\bm{z}}_{i}$ contribute to the model through the training process? Accordingly, variants of re-training, designed to estimate these contributions, are accepted as gold standards. The simplest of these is leave-one-out (LOO) re-training where each instance ${\bm{z}}_{i}$ is left out of the training set in turn to assess its contribution. For FiMODA however, it is not clear how one can “go back in time” to determine contributions to the final model. Instead, we change the question to one of sensitivity: how sensitive is the given model to a training instance ${\bm{z}}_{i}$? To measure these sensitivities, we propose further training as a gold standard, in a manner analogous to how re-training is used in the TAA setting.
@@
Further training can be described generically as follows. We start from the given parameters ${\bm{\theta}}^{f}$ and train on a dataset $\mathcal{D}^{\prime}$ using an algorithm $A^{\prime}$, i.e., ${\bm{\theta}}^{f}+\Delta{\bm{\theta}}=A^{\prime}(\mathcal{D}^{\prime},{\bm{\theta}}^{f})$, where $\mathcal{D}^{\prime}$ and $A^{\prime}$ are generally different from $\mathcal{D}$ and $A$. In this work, we restrict attention to LOO further training, corresponding to LOO datasets $\mathcal{D}^{\prime}=\mathcal{D}_{-i}\triangleq\mathcal{D}\setminus\{{\bm{z}}_{i}\}$ as well as $\mathcal{D}^{\prime}=\mathcal{D}$. This restriction aligns with the gradient-based methods discussed in Section 4; more general perturbed datasets $\mathcal{D}^{\prime}$ are of course possible. As for $A^{\prime}$, our desire is for it to be representative of neural network (NN) training. We thus assume that $A^{\prime}$ seeks to minimize the empirical risk $R(\mathcal{D}^{\prime};{\bm{\theta}})\triangleq\sum_{{\bm{z}}_{i}\in\mathcal{D}^{\prime}}L({\bm{z}}_{i};{\bm{\theta}})$:
$\Delta{\bm{\theta}}(\mathcal{D}^{\prime})\underset{\approx}{\in}\operatorname*{arg\,min}_{\Delta\theta}\>R(\mathcal{D}^{\prime};{\bm{\theta}}^{f}+\Delta{\bm{\theta}}),$ (1)
@@
where $\underset{\approx}{\in}\operatorname*{arg\,min}$ means “an approximate minimizer” (note that (1) may not have a unique minimizer), the notation ${\bm{\theta}}^{f}+\Delta{\bm{\theta}}$ is meant to indicate that the optimization is initialized at ${\bm{\theta}}^{f}$ ($\Delta{\bm{\theta}}={\bm{0}}$), and $\Delta{\bm{\theta}}(\mathcal{D}^{\prime})$ is the change in parameters resulting from applying $A^{\prime}$ to $(\mathcal{D}^{\prime},{\bm{\theta}}^{f})$.
@@
We now discuss refinements to further training to address two issues inherent to typical NN training. [^3] For convex training objectives, these issues are not much of a concern (please see Appendix A).
@@
Non-convergence: The “final” parameters ${\bm{\theta}}^{f}$ are often not a stationary point of the empirical risk. Hence, as shown in Figure 1, further training on the original training set $\mathcal{D}^{\prime}=\mathcal{D}$ generally yields a non-zero change $\Delta{\bm{\theta}}(\mathcal{D})$ and a corresponding change in the evaluation function $g({\bm{z}},{\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}))$. We refer to this as the effect due to further training alone. Similarly, $\Delta{\bm{\theta}}(\mathcal{D}_{-i})$ and $g({\bm{z}},{\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}_{-i}))$ are the effects of further training without instance ${\bm{z}}_{i}$. The difference in outputs,
@@
Stochasticity of training: The training algorithm $A^{\prime}$ also depends on random elements, notably the order of instances in each training epoch. Denoting these random elements as $\xi$, the change in parameters becomes a function of $\xi$, ${\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}^{\prime},\xi)=A^{\prime}(\mathcal{D}^{\prime},{\bm{\theta}}^{f},\xi)$, as does the evaluation function $g({\bm{z}},{\bm{\theta}}^{f}+\Delta{\bm{\theta}}(\mathcal{D}^{\prime},\xi))$. We view $\xi$ as a nuisance parameter, an artifact of the stochastic nature of modern training algorithms. Thus, we would ideally like to take an expectation over $\xi$. Applying this to the difference in (2) yields
@@
As mentioned above, further training is a natural analogue of re-training from scratch, and as shown in the next section, it is approximated by several gradient-based TDA methods. To further support the validity of further training, Appendix E.2 shows that the sensitivities quantified by further training can also detect mislabelled examples and provide insights into model behavior.